What is the de Broglie wavelength (in kilometers) of an electron (mass = 9.109e-31 kg) traveling with a velocity of 5.386e+7 m/s?
1
Expert's answer
2020-12-21T03:30:39-0500

For wavelength to be in km, the units of velocity should be turned into km/s:

v=5.386•104km/s.

L(here the Greek letter lambda should be)=h/p =h/mv =6.626•10-34/9.209•10-31×5.386•104=1.351•10-8(km).